# IMU: Influence-guided Machine Unlearning

### v2

@@ 6.2 Divergence analysis
$\|{\bm{\theta}}^{t}_{\texttt{IMU}}$- ${\bm{\theta}}^{0}\|_{{\bm{X}}^{\top}{\bm{X}}}^{2}=(\hat{{\mathcal{I}}^{c}}{\bm{a}})^{\top}{\bm{G}}(\hat{{\mathcal{I}}^{c}}{\bm{a}}),$ (21)
where $\hat{{\mathcal{I}}^{c}}=\operatorname{diag}\left(\hat{{\mathcal{I}}^{c}}({\bm{x}}_{1},{\bm{X}}),\cdots,\hat{{\mathcal{I}}^{c}}({\bm{x}}_{n_{f}},{\bm{X}})\right)$ and $\hat{{\mathcal{I}}^{c}}({\bm{x}}_{i},{\bm{X}})\in[0,1]$.
From Equations 17, 19 and 21, we can bound the divergence for GA, NPO, and IMU as
$\lambda_{\min}({\bm{G}})\cdot\|{\bm{a}}\|^{2}\leq$ $\|{\bm{\theta}}^{t}_{\texttt{GA}}-{\bm{\theta}}^{0}\|_{{\bm{X}}^{\top}{\bm{X}}}^{2}\leq\lambda_{\max}({\bm{G}})\cdot\|{\bm{a}}\|^{2},$ (22)
$\lambda_{\min}({\bm{G}})\cdot\|{\mathcal{W}}{\bm{a}}\|^{2}\leq$ $\|{\bm{\theta}}^{t}_{\texttt{NPO}}-{\bm{\theta}}^{0}\|_{{\bm{X}}^{\top}{\bm{X}}}^{2}\leq\lambda_{\max}({\bm{G}})\cdot\|{\mathcal{W}}{\bm{a}}\|^{2},$ (23)
$\lambda_{\min}({\bm{G}})\cdot\|\hat{{\mathcal{I}}^{c}}{\bm{a}}\|^{2}\leq$ $\|{\bm{\theta}}^{t}_{\texttt{IMU}}-{\bm{\theta}}^{0}\|_{{\bm{X}}^{\top}{\bm{X}}}^{2}\leq\lambda_{\max}({\bm{G}})\cdot\|\hat{{\mathcal{I}}^{c}}{\bm{a}}\|^{2},$ (24)
where $\lambda_{\min}({\bm{G}})$ and $\lambda_{\max}({\bm{G}})$ are the smallest and largest eigenvalue of the Gram matrix $\bm{G}$}, respectively. Since ${\mathcal{W}}\in[0,2]$ (typically ${\mathcal{W}}\leq 1$) and $\hat{{\mathcal{I}}^{c}}\in[0,1]$, we usually have $\|{\mathcal{W}}{\bm{a}}\|^{2}\leq\|{\bm{a}}\|^{2}$, $\|\hat{{\mathcal{I}}^{c}}{\bm{a}}\|^{2}\leq\|{\bm{a}}\|^{2}$. GA treats each sample equally and exhibits the fastest divergence, which may cause over-unlearning. NPO slows divergence by dampening high-confidence samples, while IMU reduces divergence by selectively updating only the most influential samples, promoting stability.
6.3 Derivation of influence function
The following derivation is adapted from the appendix of (Koh and Liang 2017), where the influence function is analyzed via a first-order perturbation approach. We consider the empirical risk minimizer $\bm{\theta}$$^{\ast}$, which is defined as
$R({\bm{\theta}})\coloneqq\frac{1}{n}\sum_{i=1}^{n}\ell({\bm{x}}_{i},{\bm{\theta}}),$ (25)
where $\ell({\bm{x}}_{i},{\bm{\theta}})$ denotes the loss on training sample ${\bm{x}}_{i}$. Then, assume that $R$ is twice differentiable and strongly convex w.r.t. $\bm{\theta}$}. In particular, the Hessian of the empirical risk at the optimum $\bm{\theta}$$^{\ast}$ is given as
${\bm{H}}_{{\bm{\theta}}^{\ast}}\coloneqq\nabla_{{\bm{\theta}}}^{2}R({\bm{\theta}}^{\ast})=\frac{1}{n}\sum_{i=1}^{n}\nabla_{{\bm{\theta}}}^{2}\ell({\bm{x}}_{i},{\bm{\theta}}^{\ast}),$ (26)
@@ 6.3 Derivation of influence function
${\bm{\theta}}^{\ast}_{\epsilon,{\bm{x}}}=\arg\min_{{\bm{\theta}}\in\Theta}\left\{R({\bm{\theta}})+\epsilon\ell({\bm{x}},{\bm{\theta}})\right\}.$ (27)
Then, define the parameter change under this perturbation as $\Delta_{\epsilon}={\bm{\theta}}^{\ast}_{\epsilon,{\bm{x}}}-{\bm{\theta}}^{\ast}.$ Since $\bm{\theta}$$^{\ast}$ is independent of $\epsilon$, we have $\frac{d{\bm{\theta}}^{\ast}_{\epsilon,{\bm{x}}}}{d\epsilon}=\frac{d\Delta_{\epsilon}}{d\epsilon}.$ Because ${\bm{\theta}}^{\ast}_{\epsilon,{\bm{x}}}$ minimizes the perturbed objective, it satisfies the first-order optimality condition as follows
$0=\nabla_{{\bm{\theta}}}R({\bm{\theta}}^{\ast}_{\epsilon,{\bm{x}}})+\epsilon\nabla_{{\bm{\theta}}}\ell({\bm{x}},{\bm{\theta}}^{\ast}_{\epsilon,{\bm{x}}}).$ (28)
@@
$0$ $\approx\nabla_{{\bm{\theta}}}R({\bm{\theta}}^{\ast})+\epsilon\nabla_{{\bm{\theta}}}\ell({\bm{x}},{\bm{\theta}}^{\ast})+\left[\nabla_{{\bm{\theta}}}^{2}R({\bm{\theta}}^{\ast})+\epsilon\nabla_{{\bm{\theta}}}^{2}\ell({\bm{x}},{\bm{\theta}}^{\ast})\right]\Delta_{\epsilon}.$ (29)
@@
$\Delta_{\epsilon}\approx$ $-\left[\nabla_{{\bm{\theta}}}^{2}R({\bm{\theta}}^{\ast})+\epsilon\nabla_{{\bm{\theta}}}^{2}\ell({\bm{x}},{\bm{\theta}}^{\ast})\right]^{-1}\left[\nabla_{{\bm{\theta}}}R({\bm{\theta}}^{\ast})+\epsilon\nabla_{{\bm{\theta}}}\ell({\bm{x}},{\bm{\theta}}^{\ast})\right].$ (30)
@@
$\Delta_{\epsilon}\approx-{\bm{H}}_{{\bm{\theta}}^{\ast}}^{-1}\nabla_{{\bm{\theta}}}\ell({\bm{x}},{\bm{\theta}}^{\ast})\epsilon.$ (31)
@@
$\left.\frac{d{\bm{\theta}}^{\ast}_{\epsilon,{\bm{x}}}}{d\epsilon}\right|_{\epsilon=0}=-{\bm{H}}_{{\bm{\theta}}^{\ast}}^{-1}\nabla_{{\bm{\theta}}}\ell({\bm{x}},{\bm{\theta}}^{\ast}).$ (32)
Lastly, we can define the influence function as below
${\mathcal{I}}({\bm{x}})\coloneqq-{\bm{H}}_{{\bm{\theta}}^{\ast}}^{-1}\nabla_{{\bm{\theta}}}\ell({\bm{x}},{\bm{\theta}}^{\ast}).$ (33)
7 Additional results
7.1 Results on sample-wise unlearning
In Table 7, we evaluate the unlearning performance of IMU method on ResNet-18 for sample-wise unlearning on CIFAR-10. The results clearly manifest that our proposed approach achieves robust balance across these metrics with feasible run time, compared with other existing methods. Additionally, we also examine how different influence update frequencies affect sample-wise unlearning performance in Table 6. And we can also develop the conclusion that in this scenario, infrequent updates are sufficient for effective unlearning, which validates the robustness of our method.
@@ 7.1 Results on sample-wise unlearning
Table 6: Impact of frequency ($\nu$) of influence value updates on unlearning performance. $\nu=0$ means only update at the first epoch; $\nu=1$ and $\nu=2$ mean update every epoch and every two epochs, respectively.
@@
Table 7: Quantitative results on CIFAR-10 and CIFAR-100. Performance is averaged over 10 independent runs with different random seeds for sample-wise unlearning.
7.2 Results on LLM unlearning
In Table 8 and Figure 4, we generalize our MU method on large language model and perform forgetting forget05 task on the TOFU benchmark. Compared with other LLM unlearning baselines, IMU (ours) strikes a balance on forget quality and model utility. From Figure 4, we can clearly see the catastrophic collapse issue in the GA method and slightly better performance of the NPO. Although the model structure of Llama-3.2-3B-Instruct is more complex than that of image task, which is not conducive to the accurate estimation of influence function, our method has still achieved an almost equivalent forgetting effect to SimNPO method.
Figure 4: Comparison of forget quality and model utility between GA, NPO, SimNPO, and IMU.
In Table 10, We presented $5$ randomly selected question-answer pairs from forget05 of TOFU benchmark and the corresponding output answers for different unlearning methods. Among them, the GA method always chooses not to answer or simply repeats the questions for difficult questions. Even in a simple question Q3, it did not forget the key information “Global Health”. In contrast, although NPO has also produces some meaningless information, in general, it ensured the fluency and logic of the output language. While the SimNPO method performed well in forgetting important information, but it still retained related information similar to the ground truth in some responses. However, our method IMU, despite significantly forgetting all relevant information, still maintained effective output, which greatly enhancing fluency and diversity of the generated content.
In Table 11, the question-answer pairs were selected from retain95. It means that the output of the model should be approximate to the ground truth, which proves the retention performance of the unlearned model. From Table 11, it is clearly observable that the answer provided by GA is always very simple and lacks detailed descriptions. Although NPO and SimNO’s answers are relatively accurate, they sometimes also output information that deviates from the ground truth. In contrast, our method is the closest to the facts, demonstrating the best model utility among these unlearn methods.
Method Model Utility ($\uparrow$) Extraction Strength ($\downarrow$) Forget Q A Prob ($\downarrow$) Forget Q A ROUGE ($\downarrow$) Privleak ($\downarrow$)
GA 0.00 0.03 0.03 0.00 048
NPO 0.28 0.05 0.06 0.21 080
SimNPO 0.35 0.05 0.17 0.33 -12
IMU (Ours) 0.33 0.06 0.09 0.29 -83
Table 8: Results of LLM unlearning on TOFU (Forget05) over 66 epochs.
7.3 Experiments on SalUn
Under the same dataset and experimental settings, we replicated and evaluated the current SOTA method based on the retain dataset and forget dataset SalUn (Fan et al. 2024b). The experimental results are shown in Table 9. On the CIFAR-10 dataset, the result of IMU method we proposed performs is similarly to SalUn, while on the CIFAR-100 dataset, SalUn still achieves relatively better performance.
Setting $\texttt{Acc}_{{\mathcal{D}}_{f}}(\downarrow)$ $\texttt{Acc}_{{\mathcal{D}}_{r}}(\uparrow)$ $\texttt{Acc}_{{\mathcal{D}}_{t}}(\uparrow)$ MIA $(\uparrow)$ Run time $(\downarrow)$

### v3

@@ A.2. Divergence analysis
(38) $\|{\bm{\theta}}^{t}_{\texttt{IMU}}-{\bm{\theta}}^{0}\|_{{\bm{X}}^{\top}{\bm{X}}}^{2}\leq\eta^{2}\lambda_{\max}({\bm{K}})^{2}\,\|{\bm{W}}_{\texttt{IMU}}{\bm{s}}_{\texttt{IMU}}^{(t)}\|^{2}.$
Then, for logistic regression, $\pi_{{\bm{\theta}}}(y_{i}\mid{\bm{x}}_{i})\in(0,1)$, hence $\big|(2y_{i}-1)\big(1-\pi_{{\bm{\theta}}}(y_{i}\mid{\bm{x}}_{i})\big)\big|\leq 1$. For any method $m\in\{\texttt{GA},\texttt{NPO},\texttt{IMU}\}$, $|s^{(t)}_{m,i}|\leq\sum_{k=0}^{t-1}{\bm{1}}[i_{k}=i]\leq t$. Consequently, $\|{\bm{s}}_{m}^{(t)}\|_{2}\leq\sqrt{n}\,t$ and $\|{\bm{W}}_{m}{\bm{s}}_{m}^{(t)}\|_{2}\leq\|{\bm{W}}_{m}\|_{2}\|{\bm{s}}_{m}^{(t)}\|_{2}$, we then obtain
$\|{\bm{\theta}}^{t}_{\texttt{GA}}-{\bm{\theta}}^{0}\|_{{\bm{X}}^{\top}{\bm{X}}}^{2}\leq\eta^{2}\lambda_{\max}({\bm{K}})^{2}nt^{2},$
$\|{\bm{\theta}}^{t}_{\texttt{NPO}}-{\bm{\theta}}^{0}\|_{{\bm{X}}^{\top}{\bm{X}}}^{2}\leq\eta^{2}\lambda_{\max}({\bm{K}})^{2}nt^{2}\|{\bm{W}}_{\texttt{NPO}}\|^{2},$
(39) $\|{\bm{\theta}}^{t}_{\texttt{IMU}}-{\bm{\theta}}^{0}\|_{{\bm{X}}^{\top}{\bm{X}}}^{2}\leq\eta^{2}\lambda_{\max}({\bm{K}})^{2}nt^{2}\|{\bm{W}}_{\texttt{IMU}}\|^{2}.$
GA corresponds to uniform weighting and treats each sample equally, while NPO and IMU introduce per-sample weights, which modulate the magnitude of the update directions and thus control the parameter drift.
A.3. Derivation of influence function
The following derivation is adapted from the appendix of (Koh and Liang, 2017), where the influence function is analyzed via a first-order perturbation approach. We consider the empirical risk minimizer $\bm{\theta}$$^{\ast}$, which is defined as
(40) $R({\bm{\theta}})\coloneqq\frac{1}{N}\sum_{i=1}^{N}\ell\big(f_{{\bm{\theta}}}({\bm{x}}_{i})\big),$
where $\ell(f_{{\bm{\theta}}}({\bm{x}}_{i}))$ denotes the loss on training sample ${\bm{x}}_{i}$. Then, assume that $R$ is twice differentiable and strongly convex w.r.t. $\bm{\theta}$}. In particular, the Hessian of the empirical risk at the optimum $\bm{\theta}$$^{\ast}$ is given as
(41) ${\bm{H}}_{{\bm{\theta}}^{\ast}}\coloneqq\nabla_{{\bm{\theta}}}^{2}R({\bm{\theta}}^{\ast})=\frac{1}{N}\sum_{i=1}^{N}\nabla_{{\bm{\theta}}}^{2}\ell\big(f_{{\bm{\theta}}^{\ast}}({\bm{x}}_{i})\big),$
@@ A.3. Derivation of influence function
(42) $\hat{{\bm{\theta}}}_{\epsilon,{\bm{x}}}=\arg\min_{{\bm{\theta}}\in\Theta}\left\{\frac{1}{N}\sum_{i=1}^{N}\ell\big(f_{{\bm{\theta}}}({\bm{x}}_{i})\big)+\epsilon\ell\big(f_{{\bm{\theta}}}({\bm{x}})\big)\right\}.$
Then, define the parameter change under this perturbation as $\Delta_{\epsilon}\coloneqq\hat{{\bm{\theta}}}_{\epsilon,{\bm{x}}}-{\bm{\theta}}^{\ast}.$ Since $\bm{\theta}$$^{\ast}$ is independent of $\epsilon$, we have $\frac{d\hat{{\bm{\theta}}}_{\epsilon,{\bm{x}}}}{d\epsilon}=\frac{d\Delta_{\epsilon}}{d\epsilon}.$ Because ${\bm{\theta}}^{\ast}_{\epsilon,{\bm{x}}}$ minimizes the perturbed objective, it satisfies the first-order optimality condition as follows
(43) $0=\nabla_{{\bm{\theta}}}R(\hat{{\bm{\theta}}}_{\epsilon,{\bm{x}}})+\epsilon\nabla_{{\bm{\theta}}}\ell\big(f_{\hat{{\bm{\theta}}}_{\epsilon,{\bm{x}}}}({\bm{x}})\big).$
@@
(44) $0\approx\;$ $\nabla_{{\bm{\theta}}}R({\bm{\theta}}^{\ast})+\epsilon\nabla_{{\bm{\theta}}}\ell\big(f_{{\bm{\theta}}^{\ast}}({\bm{x}})\big)+\left[\nabla_{{\bm{\theta}}}^{2}R({\bm{\theta}}^{\ast})+\epsilon\nabla_{{\bm{\theta}}}^{2}\ell\big(f_{{\bm{\theta}}^{\ast}}({\bm{x}})\big)\right]\Delta_{\epsilon}.$
@@
(45) $\Delta_{\epsilon}\approx-\left[\nabla_{{\bm{\theta}}}^{2}R({\bm{\theta}}^{\ast})+\epsilon\nabla_{{\bm{\theta}}}^{2}\ell\big(f_{{\bm{\theta}}^{\ast}}({\bm{x}})\big)\right]^{-1}\left[\nabla_{{\bm{\theta}}}R({\bm{\theta}}^{\ast})+\epsilon\nabla_{{\bm{\theta}}}\ell\big(f_{{\bm{\theta}}^{\ast}}({\bm{x}})\big)\right].$
@@
(46) $\Delta_{\epsilon}\approx-{\bm{H}}_{{\bm{\theta}}^{\ast}}^{-1}\nabla_{{\bm{\theta}}}\ell\big(f_{{\bm{\theta}}^{\ast}}({\bm{x}})\big)\,\epsilon.$
@@
(47) $\left.\frac{d\hat{{\bm{\theta}}}_{\epsilon,{\bm{x}}}}{d\epsilon}\right|_{\epsilon=0}=-{\bm{H}}_{{\bm{\theta}}^{\ast}}^{-1}\nabla_{{\bm{\theta}}}\ell\big(f_{{\bm{\theta}}^{\ast}}({\bm{x}})\big).$
Lastly, we can define the influence function as below,
(48) ${\mathcal{I}}_{\texttt{up},{\bm{\theta}}^{\ast}}({\bm{x}})\coloneqq-{\bm{H}}_{{\bm{\theta}}^{\ast}}^{-1}\nabla_{{\bm{\theta}}}\ell\big(f_{{\bm{\theta}}^{\ast}}({\bm{x}})\big).$
Appendix B Additional results
B.1. Implementation details
All the experiments were performed on NVIDIA RTX3090 with Intel Xeon processors. For image classification scenario, the original model and retrained model were both trained with over $182$ epochs using the SGD optimizer with a cosine-scheduled, and learning rate was initialized at $0.1$. For both NPO and RL, training spans $10$ epochs within the interval $10^{-4},10^{-1}]$, and $\beta$ was selected from the range $[0.1,5]$. GA’s training settings around a $5$-epoch learning rate search within the interval $10^{-5},10^{-2}]$. In the case of IU, we explored the parameter $\alpha$ within the range $[1,20]$. For SSD, searching for the parameter selection weighting $\alpha$ and dampening constant $\gamma$ was executed within the range $[1,100]$ and $[0.1,1]$, respectively, while searching for the learning rate within the range $10^{-4},10^{-2}]$. For the SCAR method, the value of temperature was defined within the range $[2,5]$, the parameter $\lambda_{2}$ was set as $0.01$, and $\lambda_{1}$ was in the range of $[1,10]$ with the learning rate searching in the interval $10^{-6},10^{-2}]$. Learning rate was selected in the range $10^{-4},10^{-2}]$ for SalUn and let sparsity ratio equal to $0.5$. We trained IMU for $5$ epochs in all experiments in this task, searching for learning rates in the range $10^{-5},10^{-2}]$ and $l_{1}$ regularization intensity in the range $2\times 10^{-4},2\times 10^{-2}]$.
For person re-identification unlearning task, original model was trained over $60$ epochs with learning rate as $3\times 10^{-4}$. In the case of IMU, it is trained for $13$ iterations with a learning rate of $1.5\times 10^{-3}$, with $\alpha$ set to $0.02$. For GA, RL, SSD, NPO, and SCAR, they are trained for $30$, $15$, 11, $25$, and $5$ epochs with learning rate searched in $1\times 10^{-5},0.1]$.
For sequence modeling unlearning problem, the original model and retrained model were both trained with over $5$ epochs using softmax activation, and the learning rate was initialized at $5\times 10^{-4}$. As for SimNPO and NPO method, the value of $\beta$ was searched in interval of $[0.1,5]$ with learning rate in range $2\times 10^{-4},8\times 10^{-3}]$, while the learning rate of GA was selected from $2\times 10^{-4},10^{-2}]$. For our IMU method, learning rate was selected as $5\times 10^{-4}$. Additionally, all the unlearning methods were conducted over $50$ iterations.
In LLM unlearning scenario, $\beta$ was selected from $[1,5]$ for NPO and SimNPO, and the learning rate was searched in the range of $10^{-6},10^{-4}]$. For IMU and GA, the learning rate was selected with $10^{-6},10^{-3}]$. All methods in this task were executed over 10 epochs and distributed across 2 NVIDIA RTX 3090 GPUs.
B.2. Extra ablation results
We conduct ablation studies to examine the effect of different design choices in the proposed unlearning framework.
Effect of Smoothing Strategies: We first evaluate several smoothing strategies applied to $w_{i}$, including sqrt, softmax, and tanh. As shown in Table 5, all methods yield comparable results, with sqrt slightly outperforming the others in both forgetting quality and model utility.
Effect of Unfreeze Strategies: When estimating influence values on the last classifier layer, updating only this layer in ResNet-18 achieves the best overall performance. Table 6 shows that unfreezing earlier layers leads to a noticeable decline in model utility, and the performance deteriorates rapidly when all layers are updated.
[…]
Effect of reweighting strategies: In Table 7, different reweighting strategies are implemented with other settings the same as IMU. Although gradient norm and sample loss can partially reflect the impact of each forget samples on the model, they are inherently local and fail to account for interactions among samples. In contrast, influence values capture how removing one sample affects the overall training objective through its coupling with other data points. Our results suggest that reweighting based on influence values leads to more effective unlearning performance.
Effect of dataset used for computing Hessian matrix: As shown in Table 8, we compare computing influence values using ${\bm{H}}^{train}$ when ${\mathcal{D}}_{r}$ is available, against ${\bm{H}}^{f}$. Utilizing the full training set to compute influence values leads to a slight improvement over the forget-only setting. This result indicates that IMU remains effective and efficient when relying only on the forget dataset.
Table 5. Effect of different smoothing strategies for processing influence values on unlearning performance.
Smooth methods $\texttt{Acc}_{{\mathcal{D}}_{f}}(\downarrow)$ $\texttt{Acc}_{{\mathcal{D}}_{r}}(\uparrow)$ $\texttt{Acc}_{{\mathcal{D}}_{t}}(\uparrow)$ $W_{dist}(\downarrow)$
none 3.18 98.50 92.80 0.09
sqrt 2.13 98.59 92.85 0.09
softmax 2.97 98.48 92.72 0.09
tanh 2.36 98.59 92.86 0.09
Table 6. Effect of unfreezing different layer combinations of ResNet-18 during the unlearning process.
Unfreeze layers $\texttt{Acc}_{{\mathcal{D}}_{f}}(\downarrow)$ $\texttt{Acc}_{{\mathcal{D}}_{r}}(\uparrow)$ $\texttt{Acc}_{{\mathcal{D}}_{t}}(\uparrow)$ $W_{dist}(\downarrow)$ Run time $(\downarrow)$
only fc 0.22 98.57 92.84 0.09 65
fc&block1 0.18 95.30 89.39 0.13 58
fc&block1-2 0.13 89.10 84.05 0.18 59
fc&block1-3 0.00 82.35 78.29 0.32 57
all layers 0.53 63.48 61.62 0.39 57
Table 7. Results on different strategies of reweighting.
Reweighting methods $\texttt{Acc}_{{\mathcal{D}}_{f}}(\downarrow)$ $\texttt{Acc}_{{\mathcal{D}}_{r}}(\uparrow)$ $\texttt{Acc}_{{\mathcal{D}}_{t}}(\uparrow)$ $W_{dist}(\downarrow)$ Run time $(\downarrow)$ none (GA) 0.48 75.89 72.43 3.59 29 gradient norm 0.93 91.66 83.16 0.58 43 sample loss 0.02 94.43 87.34 0.58 35 influence value 0.02 98.62 92.89 0.07 56
Table 8. Result of IMU when ${\mathcal{D}}_{r}$ is available for computing influence value.
Hessian type $\texttt{Acc}_{{\mathcal{D}}_{f}}(\downarrow)$ $\texttt{Acc}_{{\mathcal{D}}_{r}}(\uparrow)$ $\texttt{Acc}_{{\mathcal{D}}_{t}}(\uparrow)$ $W_{dist}(\downarrow)$ Run time $(\downarrow)$ ${\bm{H}}^{f}$ 0.02 98.62 92.89 0.07 056 ${\bm{H}}^{train}$ 0.04 99.47 93.69 0.07 345
Table 9. Effect of frequency ($\nu$) of influence value updates on unlearning performance. $\nu=0$ means only update at the first epoch; $\nu=1$ and $\nu=2$ mean update every epoch and every two epochs, respectively.
@@ B.2. Extra ablation results
B.3. Results on sample-wise unlearning
In Table 10 and Table 11, we evaluate the unlearning performance of IMU method on ResNet-18 for sample-wise unlearning on CIFAR-10 with 10% and 50% of total data as forget dataset, respectively. The results clearly manifest that our proposed approach achieves robust balance across these metrics with feasible run time, compared with other existing methods. Additionally, we also examine how different influence update frequencies affect 10% sample-wise unlearning performance in Table 9. And we can develop the conclusion that in this scenario, infrequent updates are sufficient for effective unlearning, which validates the robustness of our method.
Table 10. Quantitative results on CIFAR-10 and CIFAR-100. Performance is averaged over 10 independent runs with different random seeds for 10% sample-wise unlearning.
@@ B.3. Results on sample-wise unlearning
Table 11. Quantitative results on CIFAR-10 for 50% sample-wise unlearning.
@@
B.4. Results on distributional unlearning
Aside from vision tasks, we further evaluate on a simple sequence modeling problem. Following SimNPO (Fan et al., 2024a), we construct a mixture of Markov chains with a state space of size 10, the retain distribution consists of Markov chains that transition uniformly among states $\{1,2,3\}$, while the forget distribution is a mixture of two Markov chains Forget1 (transition uniformly among $\{4,5,6\}$) and Forget2 (transition uniformly among $\{7,8,9\}$) with equal probability. Each sequence has length $T$ and is denoted as $s=(s_{1},s_{2},\cdots,s_{T})$ where $s_{t}\in\{0,1,\cdots,9\}$ for $t\in[0,T]$. A GPT-2 model is trained to learn the conditional distribution. Details can be found in §7 of SimNPO. Table 12 presents results compared to NPO and SimNPO. Performance is evaluated by the loss values over the retain and forget sets, as well as the respective KL divergence distance calculated compared with the retrained model. IMU surpasses NPO and SimNPO in both forget quality and model utility.
B.5. Results on LLM unlearning
In Table 13 and Figure 5, we generalize our MU method on large language model and perform forgetting forget05 task on the TOFU benchmark. Compared with other LLM unlearning baselines, IMU (ours) strikes a balance on forget quality and model utility with acceptable run time. From Figure 5, we can clearly see the catastrophic collapse issue in the GA method and slightly better performance of the NPO. Although the model structure of Llama-3.2-3B-Instruct is more complex than that of image task, which is not conducive to the accurate estimation of influence function, our method has still achieved an almost equivalent forgetting effect to SimNPO method.
In Table 14, We presented $5$ randomly selected question-answer pairs from forget05 of TOFU benchmark and the corresponding output answers for different unlearning methods. Among them, the GA method always chooses not to answer or simply repeats the questions for difficult questions. Even in a simple question Q3, it still remember the key information “Global Health”. In contrast, although NPO has also produces some meaningless information, in general, it ensured the fluency and logic of the output language. While the SimNPO method performed well in forgetting important information, but it still retained related information similar to the ground truth in some responses. However, our method IMU, despite significantly forgetting all relevant information, still maintained effective output, which greatly enhancing fluency and diversity of the generated content.
In Table 15, the question-answer pairs were selected from retain95. It means that the output of the model should be approximate to the ground truth, which proves the retention performance of the unlearned model. From Table 15, it is clearly observable that the answer provided by GA is always very simple and lacks detailed descriptions. Although NPO and SimNPO’s answers are relatively accurate, they sometimes also output information that deviates from the ground truth. In contrast, our method is the closest to the facts, demonstrating the best model utility among these unlearn methods.
Table 13. Results of LLM unlearning on TOFU (Forget05) over $10$ epochs.